OpenStudy (anonymous):

Solve the following equation algebraically 3x^2=12

OpenStudy (anonymous):

\[3x^2 = 12\]\[x^2 = \frac{ 12 }{ 3 }\]\[x = \sqrt{\left(\begin{matrix}\frac{ 12 }{ 3 } \\ \end{matrix}\right)}\]\[x=2\]

OpenStudy (anonymous):

thanks